3x^2-xy=0, 2y-5x=1
Answer by stanbon(75887)

3x^2-xy=0,
2y-5x=1
--------------
Solve for "y":
y = (5x+1)/2
-----
Substitute for "y" and solve for "x":
3x^2 - x[(5x+1)/2] = 3x^2 - (5x^2+x)/2 = 0
----
5x^2 + x = 0
x(5x+1) = 0
-----
x = 0 or x = -1/5
-----
Solve for "y":
y = (5x+1)/2
If x = 0, y = 1/2
If y = (-1/5), y = 0
========================
Cheers,
Stan H.
================